ABCD is a square and P, Q are the midpoints of BC, CD respectively. If AP = a
and AQ = b, find in terms of a and b, the directed line segments (i) AB, (ii) AD,
(iii) BD and (iv) AC.
1
Expert's answer
2021-01-25T03:31:28-0500

ABCD is a square: "\overrightarrow{AB}=\overrightarrow{DC}, \overrightarrow{BC}=\overrightarrow{AD}""\overrightarrow{AB}=\overrightarrow{DC}, \overrightarrow{BC}=\overrightarrow{AD}""\overrightarrow{AB}=\overrightarrow{DC}, \overrightarrow{BC}=\overrightarrow{AD}""\overrightarrow{AB}=\overrightarrow{DC}, \overrightarrow{BC}=\overrightarrow{AD}""P, Q" are the midpoints of "BC, CD" respectively: "\overrightarrow{BP}=\dfrac{1}{2}\overrightarrow{BC}, \overrightarrow{DQ}=\dfrac{1}{2}\overrightarrow{DC}""\overrightarrow{BP}=\dfrac{1}{2}\overrightarrow{BC}, \overrightarrow{DQ}=\dfrac{1}{2}\overrightarrow{DC}""\overrightarrow{BP}=\dfrac{1}{2}\overrightarrow{BC}, \overrightarrow{DQ}=\dfrac{1}{2}\overrightarrow{DC}""\overrightarrow{BP}=\dfrac{1}{2}\overrightarrow{BC}, \overrightarrow{DQ}=\dfrac{1}{2}\overrightarrow{DC}"

"\overrightarrow{AP}=\overrightarrow{AB}+\overrightarrow{BP}=\overrightarrow{AB}+\dfrac{1}{2}\overrightarrow{BC}""\overrightarrow{AP}=\overrightarrow{AB}+\overrightarrow{BP}=\overrightarrow{AB}+\dfrac{1}{2}\overrightarrow{BC}""\overrightarrow{AP}=\overrightarrow{AB}+\overrightarrow{BP}=\overrightarrow{AB}+\dfrac{1}{2}\overrightarrow{BC}""\overrightarrow{AP}=\overrightarrow{AB}+\overrightarrow{BP}=\overrightarrow{AB}+\dfrac{1}{2}\overrightarrow{BC}""\overrightarrow{AP}=\overrightarrow{AB}+\overrightarrow{BP}=\overrightarrow{AB}+\dfrac{1}{2}\overrightarrow{BC}""\overrightarrow{AQ}=\overrightarrow{AD}+\overrightarrow{DQ}=\overrightarrow{AD}+\dfrac{1}{2}\overrightarrow{DC}""\overrightarrow{AQ}=\overrightarrow{AD}+\overrightarrow{DQ}=\overrightarrow{AD}+\dfrac{1}{2}\overrightarrow{DC}""\overrightarrow{AQ}=\overrightarrow{AD}+\overrightarrow{DQ}=\overrightarrow{AD}+\dfrac{1}{2}\overrightarrow{DC}""\overrightarrow{AQ}=\overrightarrow{AD}+\overrightarrow{DQ}=\overrightarrow{AD}+\dfrac{1}{2}\overrightarrow{DC}""\overrightarrow{AQ}=\overrightarrow{AD}+\overrightarrow{DQ}=\overrightarrow{AD}+\dfrac{1}{2}\overrightarrow{DC}"

If "\overrightarrow{AP}=\vec{a}""\overrightarrow{AP}=\vec{a}" and "\overrightarrow{AQ}=\vec{b}""\overrightarrow{AQ}=\vec{b}"

"\overrightarrow{AB}+\dfrac{1}{2}\overrightarrow{AD}=\vec{a}""\overrightarrow{AB}+\dfrac{1}{2}\overrightarrow{AD}=\vec{a}""\overrightarrow{AB}+\dfrac{1}{2}\overrightarrow{AD}=\vec{a}""\overrightarrow{AD}+\dfrac{1}{2}\overrightarrow{AB}=\vec{b}""\overrightarrow{AD}+\dfrac{1}{2}\overrightarrow{AB}=\vec{b}""\overrightarrow{AD}+\dfrac{1}{2}\overrightarrow{AB}=\vec{b}"



(i)



"\overrightarrow{AD}=\vec{b}-\dfrac{1}{2}\overrightarrow{AB}""\overrightarrow{AD}=\vec{b}-\dfrac{1}{2}\overrightarrow{AB}""\overrightarrow{AD}=\vec{b}-\dfrac{1}{2}\overrightarrow{AB}""\overrightarrow{AB}+\dfrac{1}{2}(\vec{b}-\dfrac{1}{2}\overrightarrow{AB})=\vec{a}""\overrightarrow{AB}+\dfrac{1}{2}(\vec{b}-\dfrac{1}{2}\overrightarrow{AB})=\vec{a}""\overrightarrow{AB}+\dfrac{1}{2}(\vec{b}-\dfrac{1}{2}\overrightarrow{AB})=\vec{a}""\overrightarrow{AB}+\dfrac{1}{2}(\vec{b}-\dfrac{1}{2}\overrightarrow{AB})=\vec{a}""\overrightarrow{AB}=\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b}""\overrightarrow{AB}=\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b}""\overrightarrow{AB}=\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b}"


(ii)



"\overrightarrow{AD}=\vec{b}-\dfrac{1}{2}(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{AD}=\vec{b}-\dfrac{1}{2}(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{AD}=\vec{b}-\dfrac{1}{2}(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{AD}=\vec{b}-\dfrac{1}{2}(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{AD}=\vec{b}-\dfrac{1}{2}(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{AD}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}""\overrightarrow{AD}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}""\overrightarrow{AD}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}"


(iii)



"\overrightarrow{AD}-\overrightarrow{AB}=\overrightarrow{BD}""\overrightarrow{AD}-\overrightarrow{AB}=\overrightarrow{BD}""\overrightarrow{AD}-\overrightarrow{AB}=\overrightarrow{BD}""\overrightarrow{BD}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}-(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{BD}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}-(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{BD}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}-(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{BD}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}-(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{BD}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}-(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{BD}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}-(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{BD}=2\vec{b}-2\vec{a}""\overrightarrow{BD}=2\vec{b}-2\vec{a}""\overrightarrow{BD}=2\vec{b}-2\vec{a}"


(iv)



"\overrightarrow{AD}+\overrightarrow{AB}=\overrightarrow{AC}""\overrightarrow{AD}+\overrightarrow{AB}=\overrightarrow{AC}""\overrightarrow{AD}+\overrightarrow{AB}=\overrightarrow{AC}""\overrightarrow{AC}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}+(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{AC}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}+(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{AC}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}+(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{AC}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}+(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{AC}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}+(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{AC}=\dfrac{4}{3}\vec{b}-\dfrac{2}{3}\vec{a}+(\dfrac{4}{3}\vec{a}-\dfrac{2}{3}\vec{b})""\overrightarrow{AC}=\dfrac{2}{3}\vec{a}+\dfrac{2}{3}\vec{b}""\overrightarrow{AC}=\dfrac{2}{3}\vec{a}+\dfrac{2}{3}\vec{b}""\overrightarrow{AC}=\dfrac{2}{3}\vec{a}+\dfrac{2}{3}\vec{b}"